If C=400+0.8Yd Yd=Y-T T=300+0.2Y. Find MPC
1
Expert's answer
2021-09-17T12:05:58-0400

"Given -\\
C = 400 + 0.8Yd\\ 

Yd = Y - T \\

T = 300 + 0.2Y"

Putting Value of T from above, 

"Yd = Y - (300 + 0.2Y)\\

Yd = Y - 300 - 0.2Y \\

Yd = 0.8Y - 300"

Putting value of Yd in C, 

"C = 400 + 0.8(0.8Y - 300) \\

C = 400 + 0.64Y - 240 \\

C = 160 + 0.64Y"

To Calculate MPC we find slope of Consumption function -

"MPC = \frac{∆C}{∆Y} = 0.64"